# Exploring Ground and Excited States via Single Reference Coupled-Cluster Theory and Algebraic Geometry

### v1

@@ 4.2 Parametric homotopy continuation
As $\lambda\to 0$, we may encounter different scenarios when getting close to $\lambda=0$. Here, the endgame 80 should be employed to determine the target solutions and their “type”. We illustrate such different scenarios in Figure 3.
Figure 3: Sketch of possible homotopy paths. The solid line shows a path with no finite limit as $\lambda\to 0$, the dashed lines have the same limit of a singular root, and the dotted-dashed line has a unique limit of a nonsingular root.
We conclude this section by highlighting the progress that has been made within homotopy continuation based software development exploiting parallel implementations, which can significantly extend their application areas in the coming years. In particular, we here want to emphasize PHCpack 81, Bertini, 82 HOM4PS, 83, 84 NAG4M2 74 and HomotopyContinuation.jl72
@@ 5 Numerical results
In this section we investigate the root structure and potential energy curves (PECs) for ground and excited states at the level of CCD and CCSD. We report PECs corresponding to different CC solutions obtained by the above-outlined computational procedure. The obtained energies and corresponding eigenstate approximations are then compared to exact eigenpairs of the respective Hamiltonian. Besides the computation and analysis of the CC states, we compare different bounds to the number of CC roots with the computed number of CC solutions along different bond stretching processes. We restrict our investigations to four electron systems, such as (H2)2 in D2h and D∞h configurations, H4 symmetrically disturbed on a circle 85, and LiH. The reported computations are performed in first quantization and the considered systems are all at half-filling, i.e. four electrons in eight spin orbitals. The one-body and two-body integrals for the considered systems are obtained from the Python-based Simulations of Chemistry Framework (PySCF86, 87, 88. For the algebraic simulations, we used the monodromy and parametric homotopy continuation methods in HomotopyContinuation.jl. 72
@@ 5.2 Dissociation of lithium hydride – CCD
We initiate our energy and state approximation investigations by looking at the dissociation process of lithium hydride. We consider bond distances ranging from $1.375$ to $5.95$ bohr. Diagonalizing the Hamiltonian yields the full spectrum reported in Figure 16(a) (see supplementary material) together with all roots to the CCD equations. Due to the large number of FCI solutions, we minimize the distance between FCI solutions and CCD solutions showing a much clearer picture of potentially well-resolved PECs, see Figure 16(b) in the supplementary material. This graph allows us to extract the PECs of the different states that are well-approximated using CCD, see Figure 4(a). We find that 10 PECs are approximated up to $5\cdot 10^{-2}$ hartree for the entire dissociation process, however, we also note that for various bond distances, more CCD energies approximate FCI energies well, see Figure 16(b) and Table 1. For the 10 PECs, we moreover compute the overlap between the CCD states and the approximated eigenstates, see Figure 4(b). Remarkably, three of the four lowest energy states are well approximated both in terms of energy and states. The highest energy state stands out as well since it is well-approximated near the equilibrium, both in terms of energy as well as the eigenstate. However, as we transition away from the equilibrium the level of approximation deteriorates. Note that poor overlap may be detected when the eigenspace is higher dimensional, that is if the corresponding eigenvalue has multiplicity two or higher. Therefore some of the higher energy states reporting bad overlap might have CCD states that are close to the actual eigenspace – or even within the eigenspace – since we are only reporting overlap with one representative from the eigenspace. A detailed investigation of this is left for future works.
@@
Figure 4: (a) PECs of LiH that can be accurately approximated by CCD energies. (b) Overlaps of the CCD states with the corresponding eigenstate.
@@
Investigating the root structure, we observe that the total number of CCD roots fluctuates along the bond stretching process, see Table 1. We find that any bound for a generic Hamiltonian severely overestimates the true number of roots for this system.
@@ 5.4 Dissociating systems of hydrogen
We now investigate the dissociation of (H2)2 planar model systems in different geometries 38.
@@
The full spectrum together with the CCD solutions can be found in Figure 13(a) in the Supplementary material, and the PECs together with the closest CCD solutions are reported in Figure 13(b). We extract the PECs that are well-approximated using CCD and report them in Figure 8(a). Here, we observe that eight CCD energies approximate FCI PECs well. For these eight CCD states, we compute the overlap with the corresponding eigenstate, see Figure 8(b). We see that the ground state is well approximated, both in terms of energy as well as the state. The following six states (i.e. states two to seven in Figure 8(b)) have intermediate overlaps with the targeted states; the last state has a poor overlap with the targeted state. This illustrates that CC amplitudes could yield good energy approximations while providing poor approximations to the actual eigenstates 13.
Figure 8: (a) PECs of (H2)2 in D2h symmetry that are accurately described by CCD energies. (b) Overlap of the CCD states with the corresponding eigenstate.
@@
The full spectrum and CCD solutions is presented in Figure 14(a) in the Supplementary material, with the PECs and their closest CCD solutions detailed in Figure 14(b). We have identified PECs that are accurately approximated by CCD, as shown in Figure 10(a), where four CCD energies closely match the FCI PECs. The overlap of these four CCD states with their respective eigenstates is analyzed in Figure 10(b). Consistent with observations from the (H2)2 in D2h configuration, the ground state demonstrates a high degree of accuracy, while the other three states exhibit intermediate to poor overlap with their targeted eigenstates.
Figure 10: (a) PECs of (H2)2 in D∞h symmetry that are accurately described by CCD energies. (b) Overlap of the CCD states with the corresponding eigenstate.
@@ 5.5 H4 disturbed on a circle
We proceed by investigating a variant of the H4 model consisting of four hydrogen atoms symmetrically distributed on a circle of radius $R=\sqrt{2}$ bohr 89, see Fig. 11.
@@
The full spectrum together with the CCD solutions can be found in Figure 15(a) in the Supplementary material, and the PECs together with the closest CCD solutions are reported in Figure 15(b). We extract the PECs that are well-approximated using CCD which are reported in Figure 12(a). Note that it is well-known that for this system CCD does not perform well in the region close to 90, due to strong degeneracies 38. This results in a dramatic reduction of the total number of roots, we find three roots at $90^{\circ}$ and six roots at $89^{\circ}$. For completion, these energies are included in the supplementary material, however, since the focus of this work is on the root structure and potentially physically relevant roots accessible by single-reference CC theory, we focus on cases where single-reference CC theory provides reasonable approximations and therefore excluded these points in Figure 12(a). Outside of this challenging region, we find five solutions that approximate PECs well. For these five solutions, we compute the overlap with the corresponding eigenstate, see Figure 12(b). Consistent with the previous H4 systems, this model system shows the effect of CCD states accurately resolving excited energies but providing poor overlap to the exact eigenstates. The only state that shows a good approximation to its eigenstate is the CCD ground state.
Figure 12: (a) PECs of H4 symmetrically distributed on a cycle that can be accurately reconstructed by CCD roots. (b) Overlap of the CCD states with the corresponding eigenstate.
@@ 6 Conclusion
Additionally, we examined the total number of roots and the effectiveness of various bounds developed for generic Hamiltonians. Our simulations indicate that while the most recent bounds perform exceptionally well for generic systems, they significantly overestimate the number of solutions in physical systems. More precisely, for all considered systems, we computed between $30$-$60$ real solutions of which only $10$-$30$ approximate physical states, compared to CC degree of $72$. This excess of solutions gets amplified for CCSD: Solving the CC equations of a generic Hamiltonian yields 16 952 9961695299616\,952\,99616 952 996 solutions, compared to $1280$ solutions that yield real-valued energies of the CC equations describing LiH. This overestimation by a factor of over ten thousand indicates that to further improve the bounds of the CC equations, specific structures of the Hamiltonian have to be taken into account.
@@
The authors thank Piotr Piecuch for insightful discussions.

### v2

@@ 4.2 Parametric homotopy continuation
As $\lambda\to 0$, we may encounter different scenarios when getting close to $\lambda=0$. Here, the endgame 89 should be employed to determine the target solutions and their “type”. We illustrate such different scenarios in Figure 3.
Figure 3: Sketch of possible homotopy paths65. The solid line shows a path with no finite limit as $\lambda\to 0$, the dashed lines have the same limit of a singular root, and the dotted-dashed line has a unique limit of a nonsingular root.
We conclude this section by highlighting the progress that has been made within homotopy continuation based software development exploiting parallel implementations, which can significantly extend their application areas in the coming years. In particular, we here want to emphasize PHCpack 90, Bertini, 91 HOM4PS, 92, 93 NAG4M2 82 and HomotopyContinuation.jl80
@@ 5 Numerical results
In this section we investigate the root structure and potential energy curves (PECs) for ground and excited states at the level of CCD and CCSD. We report PECs corresponding to different CC solutions obtained by the above-outlined computational procedure. The obtained energies and corresponding eigenstate approximations are then compared to exact eigenpairs of the respective Hamiltonian. Besides the computation and analysis of the CC states, we compare different bounds to the number of CC roots with the computed number of CC solutions along different bond stretching processes. We restrict our investigations to four electron systems, such as (H2)2 in D2h and D∞h configurations, H4 symmetrically disturbed on a circle 94, and LiH. The reported computations use a minimal basis description using the STO6G basis set and all considered systems at half-filling, i.e. four electrons in eight spin orbitals. The one-body and two-body integrals for the considered systems are obtained from the Python-based Simulations of Chemistry Framework (PySCF95, 96, 97. For the algebraic simulations, we used the monodromy and parametric homotopy continuation methods in HomotopyContinuation.jl. 80
@@ 5.2 Dissociation of lithium hydride – CCD
We initiate our energy and state approximation investigations by looking at the dissociation process of lithium hydride. We consider bond distances ranging from $1.375$ to $5.95$ bohr. Diagonalizing the Hamiltonian yields the full spectrum reported in Figure LABEL:fig:CCD_LiH_full (see supplementary material) together with all roots to the CCD equations. Due to the large number of FCI solutions, we minimize the distance between FCI solutions and CCD solutions showing a much clearer picture of potentially well-resolved PECs, see Figure LABEL:fig:CCD_LiH_minima in the supplementary material. This graph allows us to extract the PECs of the different states that are well-approximated using CCD, see Figure 4(a). We find that 10 PECs are approximated up to $5\cdot 10^{-2}$ hartree for the entire dissociation process, however, we also note that for various bond distances, more CCD energies approximate FCI energies well, see Figure LABEL:fig:CCD_LiH_minima and Table 1. For the 10 PECs, we moreover compute the overlap between the CCD ket-states, i.e., $|{\Psi}\rangle=e^{\hat{T}}|{\Phi_{0}}\rangle$, and the approximated eigenstates, see Figure 4(b). We find that three of the four lowest energy states are well approximated both in terms of energy and states. The highest energy state stands out as well since it is well-approximated near the equilibrium, both in terms of energy as well as the eigenstate. However, as we transition away from the equilibrium the level of approximation deteriorates. A potential explanation for this observation is that the 1s orbital of the lithium atom forms a core making particle-hole excitations from the lowest orbital of LiH much less important than those from the valence $\sigma$ orbital. Thus, LiH is quite close to a two-electron problem for which CCSD is exact, and therewith, one could argue that the resulting near-exactness of CCSD for LiH means that the solutions of the CCSD equations (including wave functions) are expected to be numerically close to their full CI counterparts. We also note that a number of states show rather poor overlap with the corresponding eigenstates, which aligns with previous findings when investigating the root structure of H4 clusters13. There are several potential reasons for this. First, the poor overlap may be detected when the eigenspace is higher dimensional, that is if the corresponding eigenvalue has multiplicity two or higher. Therefore some of the higher energy states reporting bad overlap might have CCD states that are close to the actual eigenspace – or even within the eigenspace – since we are only reporting overlap with one representative from the eigenspace. Second, it might be that the poor resolution via the ket-state is related to CC being a non-variational theory. In this case, the inclusion of a projection onto the bra-state, i.e., $e^{\hat{T}}|{\Phi_{0}}\rangle\langle{\Phi_{0}}|(I+\hat{\Lambda})e^{-\hat{T}}$ where $\hat{\Lambda}$ is a de-excitation operator defined via the dual variable of $t$ in the CC Lagrange formulation, could yield a more detailed insight. A thorough investigation of this is left for future work.
@@
Figure 4: (a) PECs of LiH (solid lines) that can be accurately approximated by CCD energies (crosses). (b) Overlaps of the CCD states with the corresponding eigenstate. The dashed red line indicates the equilibrium geometry.
@@
Investigating the root structure, we observe that the total number of CCD roots fluctuates along the bond stretching process, see Table 1. This fluctuation could be caused by the continuous parameter curve of Hamiltonians crossing through the discriminant of the CC family, yielding a change in the root count. We find that any bound for a generic Hamiltonian severely overestimates the true number of roots for this system.
@@ 5.4 Dissociating systems of hydrogen
We now investigate the dissociation of (H2)2 planar model systems in different geometries 49, 98, 13.
@@
The full spectrum together with the CCD solutions can be found in Figure LABEL:fig:CCD_h4_D4_full in the Supplementary material, and the PECs together with the closest CCD solutions are reported in Figure LABEL:fig:CCD_D4_min. We extract the PECs that are well-approximated using CCD and report them in Figure 8(a). Here, we observe that eight CCD energies approximate FCI PECs well. For these eight CCD states, we compute the overlap with the corresponding eigenstate, see Figure 8(b). We see that the ground state is well approximated, both in terms of energy as well as the state. The following six states (i.e. states two to seven in Figure 8(b)) have intermediate overlaps with the targeted states; the last state has a poor overlap with the targeted state.
Figure 8: (a) PECs of (H2)2 in D2h symmetry (solid lines) that are accurately described by CCD energies (crosses). (b) Overlap of the CCD states with the corresponding eigenstate.
@@
The full spectrum and CCD solutions is presented in Figure LABEL:fig:CCD_h4_P4_full in the Supplementary material, with the PECs and their closest CCD solutions detailed in Figure LABEL:fig:CCD_h4_P4_min. We have identified PECs that are accurately approximated by CCD, as shown in Figure 10(a), where four CCD energies closely match the FCI PECs. The overlap of these four CCD states with their respective eigenstates is analyzed in Figure 10(b). Consistent with observations from the (H2)2 in D2h configuration, the ground state demonstrates a high degree of accuracy, while the other three states exhibit intermediate to poor overlap with their targeted eigenstates.
Figure 10: (a) PECs of (H2)2 in D∞h symmetry (solid lines) that are accurately described by CCD energies (crosses). (b) Overlap of the CCD states with the corresponding eigenstate.
@@ 5.5 H4 disturbed on a circle
We proceed by investigating a variant of the H4 model consisting of four hydrogen atoms symmetrically distributed on a circle of radius $R=\sqrt{2}$ bohr 99, 100, 101. We here investigate different geometries defined by the parameter $\Theta\in(0^{\circ},180^{\circ})$ which describes the angle formed by the line segments connecting two hydrogen atoms that are placed point symmetrically with respect to the origin of the circle. For a schematic depiction of this process see Fig. 11.
@@
The full spectrum together with the CCD solutions can be found in Figure LABEL:fig:CCD_h4_P4D4_full in the Supplementary material, and the PECs together with the closest CCD solutions are reported in Figure LABEL:fig:CCD_h4_P4D4_min. We extract the PECs that are well-approximated using CCD which are reported in Figure 12(a). Note that it is well-known that for this system CCD does not perform well in the region close to 90, due to strong degeneracies 49. This results in a dramatic reduction of the total number of roots, we find three roots at $90^{\circ}$ and six roots at $89^{\circ}$. For completion, these energies are included in the supplementary material, however, since the focus of this work is on the root structure and potentially physically relevant roots accessible by single-reference CC theory, we focus on cases where single-reference CC theory provides reasonable approximations and therefore excluded these points in Figure 12(a). Outside of this challenging region, we find five solutions that approximate PECs well. For these five solutions, we compute the overlap with the corresponding eigenstate, see Figure 12(b). Consistent with the previous H4 systems, this model system shows the effect of CCD states accurately resolving excited energies but providing poor overlap to the exact eigenstates. The only state that shows a good approximation to its eigenstate is the CCD ground state.
Figure 12: (a) PECs of H4 symmetrically distributed on a cycle (solid lines) that can be accurately reconstructed by CCD roots (crosses). (b) Overlap of the CCD states with the corresponding eigenstate.
@@ 6 Conclusion
Additionally, we examined the total number of roots and the effectiveness of various bounds developed for generic Hamiltonians. Our simulations indicate that while the most recent bounds perform exceptionally well for generic systems, they significantly overestimate the number of solutions in physical systems. More precisely, for all considered systems, we computed between $30$-$60$ real solutions of which only $10$-$30$ approximate physical states, compared to CC degree of $72$. This excess of solutions gets amplified for CCSD: Solving the CC equations of a generic Hamiltonian yields 16 952 9961695299616\,952\,99616 952 996 solutions, compared to $1280$ solutions that yield real-valued energies of the CC equations describing LiH. This overestimation by a factor of over ten thousand indicates that to further improve the bounds of the CC equations, specific structures of the Hamiltonian have to be taken into account. This includes e.g. a spin-integrated or spin-adapted description making the system of polynomial equations more physical and closer to conventional computational quantum chemistry formulations.
@@
The authors thank Piotr Piecuch for insightful discussions. Additionally, we thank the anonymous reviewers for their comments, which have significantly contributed to the improvement of this manuscript.
7 Associated Content
Supporting Information Available:
We present graphical representations of the full energy spectra alongside the real-valued CCD energies for several Hamiltonians: (H2)2 in both D2h and D∞h configurations, H4 symmetrically distorted on a circle, and lithium hydride. Additionally, we analyze the CCD energies by minimizing the difference between these energies and the spectral values of the Hamiltonians. This yields the CCD energies that best approximate the spectral values of the Hamiltonians.
This information is available free of charge at the website: http://pubs.acs.org/